# A Simple First-Order Algorithm for Full-Rank Equality Constrained Optimization

### v1

@@ 3.4 Telescoping sum
We now combine the effects of tangential and normal steps to accumulate in a telescoping sum the relevant decreases on $\psi(x)$ across all iterations.
@@
Suppose that AS.0–AS.7 hold. Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right),$ (3.14) where $\kappa_{\mbox{\tiny gap}}=\frac{2\sqrt{2}}{\eta}\left(\psi(x_{0})+\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}+\rho\kappa_{c}-f_{\rm low}\right).$
  • Proof.    Combining (3.2), (3.7) and (3.8), we obtain that
@@
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\left(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\right)$ $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$ (3.15)
    Then, if $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$, we have that $w_{k_{\tau_{0}}}=\varsigma$ and $\Gamma_{k_{\tau_{0}}}=0$. The bounds (3.11) and (3.15) therefore give that
    $\psi(x_{k+1})$ $-\psi(x_{0})$ (3.16)
    $=\sum_{\tau=\tau_{0}}^{\tau_{1}}\big(\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\big(\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\big)$
@@
    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta^{2}\kappa_{\mbox{\tiny tan}}\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right)$
@@
    $\leq-\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}-\frac{\eta}{2\sqrt{2}}\,\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|+\eta\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$
    where we used the fact that $\eta\leq 1$. But (2.3), AS.2, AS.4, and AS.5 ensure that
    $\|\widehat{\lambda}(x)\|\leq\frac{\kappa_{J}}{\sigma_{0}^{2}}\|g(x)\|\leq\frac{\kappa_{J}\kappa_{g}}{\sigma_{0}^{2}}.$
    Hence, using (2.5), AS.1, AS.2, AS.3, we have that
@@
    $\psi(x_{k+1})-\psi(x_{0})\geq f_{\rm low}-\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}^{2}}-\rho\kappa_{c}-\psi(x_{0})\stackrel{{\scriptstyle\rm def}}{{=}}-\eta\kappa_{\mbox{\tiny gap}},$
@@
This result is central in our analysis as it provides consistent upper bounds on a quantity related to the optimality mesure $\|g_{T,k}\|+\|c_{k}\|$, which we now exploit separately.
@@ 3.5 Tangential complexity
  • Proof.    Suppose that $t>e^{b/c}$. Then $b\leq c\,\log(t)$ and thus (see [28, eq. (14)])
@@
Suppose that AS.0–AS.7 hold. Then $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{T}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$ (3.17) and $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\leq\kappa_{T}\sqrt{\tau_{1}+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right).$ (3.18)
@@
    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right).$
@@
    $1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\leq\frac{2}{\varsigma}\Gamma_{k_{\tau_{1}+1}}$
@@
    $\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)=\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\right).$
@@
    $\sqrt{\Gamma_{k_{\tau_{1}+1}}}=\sqrt{\frac{\varsigma}{2}}\sqrt{\frac{2\Gamma_{k_{\tau_{1}+1}}}{\varsigma}}\leq\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{2\kappa_{\mbox{\tiny tan}}}},\frac{32\kappa_{\mbox{\tiny tan}}^{2}}{\varsigma}\right]$
@@
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\sqrt{\tau_{1}+1}\,\sqrt{\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|^{2}}=\sqrt{\tau_{1}+1}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}\leq\sqrt{\tau_{1}+1}\,\kappa_{T}.$ (3.19)
@@
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\|c_{k_{\tau}}\|\leq\sum_{\tau=\tau_{0}}^{\tau_{1}}\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta}{\sqrt{\varsigma}}\sum_{\tau=\tau_{0}}^{\tau_{1}}\|g_{T,k_{\tau}}\|\leq\frac{\beta\eta\kappa_{T}\sqrt{\tau_{1}+1}}{\sqrt{\varsigma}}.$
@@
Note that a marginally tighter bound on $\Gamma_{k_{\tau_{1}}}$ can obtained, as in [2], by using the Lambert $W_{-1}$ function (see [6]) instead of applying Lemma 3.6, but we have chosen the latter for simplicity of exposition.
@@ 3.6 Normal complexity
Suppose that AS.0–AS.7 hold. Then $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|<\kappa_{N}\,\left(1+\frac{2\sqrt{2}}{\beta\eta}\right),$ (3.20) where $\kappa_{N}=\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\kappa_{T}}{\varsigma}\right).$ (3.21)
@@
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(1+\frac{\Gamma_{k_{\tau_{1}+1}}}{\varsigma}\right)$ (3.22)
    where $k_{\tau_{1}}$ is the index of the last tangential iteration before $k_{\nu_{1}}$. As in Lemma 3.7, we may now use the assumption that $\Gamma_{k_{\tau_{1}}}\geq\varsigma$ and the monotonicity of the logarithm to obtain that
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right)\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2}{\varsigma}}\kappa_{T}\right)=\kappa_{N}.$ (3.23)
    Using the switching condition (2.12), we obtain that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$,
    $\|c_{k_{\nu}}\|\geq\beta\,\alpha_{T,k_{\nu}}\|g_{T,k_{\nu}}\|.$ (3.24)
    Observe now that, using the assumption that $\Gamma_{k_{\nu}}\geq\varsigma$, and (3.17),
    $\alpha_{T,k_{\nu}}=\frac{\eta}{\sqrt{\Gamma_{k_{\nu}}+\varsigma}}\geq\frac{\eta}{\sqrt{2\Gamma_{k_{\nu}}}}\geq\frac{\eta}{\sqrt{2\kappa_{T}}}.$
    Thus, (3.24) yields that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$
    $\|c_{k_{\nu}}\|\ \geq\frac{\beta\eta}{\sqrt{2\kappa_{T}}}\|g_{T,k_{\nu}}\|.$ (3.25)
@@
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\|g_{T,k_{\nu}}\|\leq\frac{\sqrt{2\kappa_{T}}}{\beta\eta}\sum_{\nu=\nu_{0}}^{\nu_{1}}\|c_{k_{\nu}}\|\leq\frac{\kappa_{N}\,\sqrt{2\kappa_{T}}}{\beta\eta}.$
    Summing this bound with (3.23) gives (3.20). $\Box$

### v2

@@ 3.4 Telescoping sum
We now combine the effects of tangential and normal steps to accumulate in a telescoping sum the relevant decreases on $\psi(x)$ across all iterations. This requires an additional assumption involving where
$\mathbb{P}_{T,j}\stackrel{{\scriptstyle\rm def}}{{=}}\mathbb{P}_{j}\!\left[\frac{\alpha_{T,j}\|g_{j}\|}{\|c_{j}\|}\right]$
(We assume that $\mathbb{P}_{T,j}=1$ if $\|c_{j}\|=0$.) Note that $\mathbb{P}_{T,j}$ is the conditional probability that iteration $j$ uses a tangential step, i.e. the conditional probability that $j\in\{k_{\tau}\}$. We assume that
AS.11:
There exists a constant $\mu_{T}\in(0,1]$ such that
$\sum_{j}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\mathbb{P}_{T,j}\right]\geq\mu_{T}\sum_{j\in{k_{\tau}}}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\right]$
@@
Suppose that AS.0–AS.10 hold. Then $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]+\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right],$ (3.16) where $\kappa_{\mbox{\tiny gap}}=\frac{1}{\eta}\left(\psi(x_{0})+\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}}-f_{\rm low}\right)+\varsigma.$
  • Proof.    Consider first the effect of tangential steps. We have that
    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(-\alpha_{T,j}\|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\right]\mathbb{P}_{T,j}\right]$
    $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(-\alpha_{T,j}\|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\mathbb{P}_{T,j}\right]$
    where we used (3.4) to derive the first inequality and the independence of the two gradient estimations to deduce the final equality. Now AS.11, (3.9) and (3.10) give that
    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ $\leq=-\mu_{T}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}\|g_{T,j}\|^{2}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
@@
    $\leq\eta\mu_{T}\,\mathbb{E}_{k}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]+\eta\mu_{T}\kappa_{\mbox{\tiny tan}}\mathbb{E}_{j}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]$
    For normal steps,
    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\nu}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right](1-\mathbb{P}_{T,j})\right]=-\eta\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\|c_{j}\|(1-\mathbb{P}_{T,j})\right]\leq-\eta\mu_{N}\sum_{j=0}^{k}\|c_{j}\|\mathbb{E}_{j}\!\left[(1-\mathbb{P}_{T,j})\right]$
    Combining (3.4), (3.9) and (3.10) and taking total expectation, we obtain that
    $\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(\psi(x_{j+1})\right]-\psi(x_{j})\right)$ $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\mathbb{P}_{T,j}\right]$ (3.17)
    $=\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\tau}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right]\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]$
    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\mathbb{E}^{\nu}_{j}\!\left[\left(\psi(x_{j+1})-\psi(x_{j})\right)\right](1-\mathbb{P}_{T,j})\right]$
    $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\left(-\alpha_{T,j}|g_{T,j}\|^{2}+\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right)\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]$
    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)(1-\mathbb{P}_{T,j})\right]$
    $\leq\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[-\alpha_{T,j}|g_{T,j}\|^{2}\right]\mathbb{E}_{j}\!\left[\mathbb{P}_{T,j}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
    $\hskip 28.45274pt+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)(1-\mathbb{P}_{T,j})\right]$
    $\leq\mu_{T}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[-\alpha_{T,j}|g_{T,j}\|^{2}\right]+\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[\alpha_{T,j}^{2}\|g_{T,j}\|^{2}\right]$
    $\hskip 28.45274pt+\mu_{N}\sum_{j=0}^{k}\mathbb{E}_{j}\!\left[(-\eta\|c_{j}\|)\right]$
@@
    $\leq-\eta\mu\,\mathbb{E}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]+???$
    Then, if $\min[k_{\nu_{0}},k_{\tau_{0}}]=0$ and $\max[k_{\nu_{1}},k_{\tau_{1}}]=k$, we have that $w_{k_{\tau_{0}-1}}=\varsigma$ and $\Gamma_{k_{\tau_{0}-1}}=0$. The bounds (3.13) and (3.17) therefore give that
    $\mathbb{E}\!\left[\psi(x_{k+1})\right]$ $-\psi(x_{0})$ (3.18)
    $=\sum_{\tau=\tau_{0}}^{\tau_{1}}\big(\mathbb{E}\!\left[\psi(x_{k_{\tau}+1})\right]-\mathbb{E}\!\left[\psi(x_{k_{\tau}})\right]\big)+\sum_{\nu=\nu_{0}}^{\nu_{1}}\big(\mathbb{E}\!\left[\psi(x_{k_{\nu}+1})\right]-\mathbb{E}\!\left[\psi(x_{k_{\nu}})\right]\big)$
@@
    $\leq-\eta\,\mathbb{E}\!\left[w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\right]-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right)\right]$
    $\leq-\eta\left(\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]-\sqrt{\varsigma}\right)-\eta\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]+\eta\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]$
    where we used AS.5. But AS.4, AS.5 and AS.7 ensure that
    $\widehat{\lambda}(x)\leq\frac{\kappa_{J}}{\sigma_{0}}\|G(x)\|\leq\frac{\kappa_{J}\kappa_{g}}{\sigma_{0}}.$
    Hence, using (2.5), AS.1, AS.2, AS.3, we have that, for all realizations,
@@
    $\psi(x_{k+1})-\psi(x_{0})\geq f_{\rm low}-\frac{\kappa_{J}\kappa_{g}\kappa_{c}}{\sigma_{0}}-\psi(x_{0})\stackrel{{\scriptstyle\rm def}}{{=}}-\eta(\kappa_{\mbox{\tiny gap}}+\varsigma),$
@@
This result is central in our analysis as it provides consistent upper bounds on a quantity related to the expected value of the optimality mesure $\|g_{T,k}\|+\|c_{k}\|$, which we now exploit separately.
@@ 3.5 Tangential complexity
  • Proof.    Suppose that $t>e^{b/c}$. Then $b\leq c\,\log(t)$ and thus (see [26, eq. (14)])
@@
Suppose that AS.0–AS.10 hold. Then $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\kappa_{T}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{\kappa_{\mbox{\tiny tan}}}},2\kappa_{\mbox{\tiny gap}}^{2}\right]$ (3.19) and $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|+\|c_{k_{\tau}}\|\right]\leq\kappa_{T}\sqrt{\tau_{1}+1}\left(1+\frac{\beta\eta}{\sqrt{\varsigma}}\right)$ (3.20)
@@
    $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right].$
@@
    $1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\leq\frac{2}{\varsigma}\Gamma_{k_{\tau_{1}}}$
@@
    $\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]=\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right)\right].$
    We may now use Jensen’s inequality and the concavity of the logarithm to deduce that
    $\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\right).$
@@
    $\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]=\sqrt{\frac{\varsigma}{2}}\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\leq\sqrt{\frac{\varsigma}{2}}\max\left[e^{\frac{\kappa_{\mbox{\tiny gap}}}{\kappa_{\mbox{\tiny tan}}}},2\kappa_{\mbox{\tiny gap}}^{2}\right]$
@@
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|g_{T,k_{\tau}}\|\right]\leq\sqrt{\tau_{1}+1}\,\mathbb{E}\!\left[\sqrt{\sum_{j=0}^{k}\|g_{T,k_{\tau}}\|^{2}}\right]=\sqrt{\tau_{1}+1}\,\mathbb{E}\!\left[\sqrt{\Gamma_{k_{\tau_{1}}}}\right]\leq\sqrt{\tau_{1}+1}\,\kappa_{T},$
    which, with (3.2), gives that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|G_{T,k_{\tau}}\|\right]\leq\kappa_{T}\sqrt{\tau_{1}+1}.$ (3.21)
@@
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|c_{k_{\tau}}\|\right]\leq\sum_{\tau=\tau_{0}}^{\tau_{1}}\beta\mathbb{E}\!\left[\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|\right]\leq\frac{\beta\eta}{\sqrt{\varsigma}}\sum_{\tau=\tau_{0}}^{\tau_{1}}\mathbb{E}\!\left[\|g_{T,k_{\tau}}\|\right]\leq\frac{\beta\eta\kappa_{T}\sqrt{\tau_{1}+1}}{\sqrt{\varsigma}}.$
@@
Note that a marginally tighter bound on $\mathbb{E}\!\left[\Gamma_{k_{\tau_{1}}}\right]$ can obtained, as in [1], by using the Lambert $W_{-1}$ function (see [5]) instead of applying Lemma 3.6, but we have chosen the latter for simplicity of exposition.
@@ 3.6 Normal complexity
Suppose that AS.0–AS.10 hold. Then $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|+\|c_{k_{\nu}}\|\right]<\frac{4\kappa_{N}\,\kappa_{g}}{\beta\eta}\sqrt{k_{\nu}+1},$ (3.22) where $\kappa_{N}=\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\log\left(\frac{2\kappa_{T}}{\varsigma}\right).$ (3.23)
@@
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+\kappa_{\mbox{\tiny tan}}\mathbb{E}\!\left[\log\left(1+\frac{\Gamma_{k_{\tau_{1}}}}{\varsigma}\right)\right]$ (3.24)
    where $k_{\tau_{1}}$ is the index of the last tangental iteration before $k_{\nu_{1}}$. As in Lemma 3.7, we may now use the assumption that $\Gamma_{k_{\tau_{1}}}\geq\varsigma$, Jensen’s inequality and the concavity and monotonicity of the logarithm to obtain that
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\mathbb{E}\!\left[\sqrt{\frac{2\Gamma_{k_{\tau_{1}}}}{\varsigma}}\right]\right)\leq\kappa_{\mbox{\tiny gap}}+2\kappa_{\mbox{\tiny tan}}\log\left(\sqrt{\frac{2}{\varsigma}}\kappa_{T}\right)=\kappa_{N}.$ (3.25)
    Taking now the conditional expectation in the switching condition (2.12), we obtain that, for $\nu\in\{\nu_{0},\ldots,\nu_{1}\}$,
    $\|c_{k_{\nu}}\|\geq\beta\,\mathbb{E}_{k_{\nu}}\!\left[\alpha_{T,k_{\nu}}\|g_{T,k_{\nu}}\|\right].$
    Observe now that, using the assumption that $\Gamma_{k_{\nu}}\geq\varsigma$, (2.10) and AS.9,
    $\alpha_{T,k_{\nu}}=\frac{\eta}{\sqrt{\Gamma_{k_{\nu}}+\varsigma}}\geq\frac{\eta}{\sqrt{2\Gamma_{k_{\nu}}}}\geq\frac{\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}.$
    Thus, using (3.2),
    $\|c_{k_{\nu}}\|\ \geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\mathbb{E}_{k_{\nu}}\!\left[\|g_{T,k_{\nu}}\|\right]\geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\|G_{T,k_{\nu}}\|,$
    in turn yielding that
    $\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\geq\frac{\beta\eta}{\kappa_{g}\sqrt{2(k_{\nu}+1)}}\mathbb{E}\!\left[\|G_{T,k_{\nu}}\|\right].$ (3.26)
@@
    $\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|G_{k_{\nu}}\|\right]\leq\frac{\kappa_{g}\sqrt{2(k_{\nu}+1)}}{\beta\eta}\sum_{\nu=\nu_{0}}^{\nu_{1}}\mathbb{E}\!\left[\|c_{k_{\nu}}\|\right]\leq\frac{\kappa_{N}\,\kappa_{g}\sqrt{2(k_{\nu}+1)}}{\beta\eta}.$
    Summing this bound with (3.25) and using $\sqrt{\kappa_{g}}\leq\beta\eta$ gives (3.22). $\Box$